# Thermal transport mapping in twisted double bilayer graphene

### v1

@@ Abstract
Two-dimensional (2D) materials have attracted significant interest due to their tunable physical properties when stacked into heterostructures. Twisting adjacent layers introduces moiré patterns that strongly influence the material’s electronic and thermal behavior. In twisted graphene systems, the twist angle critically alters phonon transport, leading to reduced thermal conductivity compared to Bernal-stacked configurations. However, experimental investigations into thermal transport in twisted structures remain limited. Here, we study the local thermal properties of twisted double bilayer graphene (TDBG) using Scanning Thermal Microscopy (SThM). We find a reduction in thermal resistance of $0.3\pm 0.1\times 10^{6}$ K W-1 compared to untwisted bilayers, attributed to changes in both intrinsic thermal conductivity and the tip–sample interface. These results, supported by analytical modeling, provide new insight into thermal transport mechanisms in twisted 2D systems and offer a pathway toward thermal engineering in twistronic devices.
In recent years, two-dimensional (2D) materials have gained increasing attention. The atomic layers in these materials are held together by weak van der Waals forces, which makes them easy to cleave and stack into new heterostructuresGeim and Grigorieva (2013). Stacking different 2D layers can result in artificial systems with novel physical propertiesSutter and Sutter (2021); Li, Zhou, and Zhang (2017); Wu et al. (2022). Moreover, by twisting the layers with respect to each other, the periodicity of the resulting lattice changes, forming a so-called moiré patternHe et al. (2021). This twist angle becomes a crucial parameter for tuning the band structureRen et al. (2020) and the lattice symmetry of the new materialCarr et al. (2017); Z. and S. (2021). At small twist anglesKerelsky et al. (2019), the mismatch between layers leads to atomic reconstruction, favoring energetically stable stacking configurations. This results in discrete stacking domains and domain wallsYoo et al. (2019), which are predicted to significantly affect the local thermal properties of twisted 2D systems.
The thermal conductivity of graphene is known to depend on the twist angle between its layers. Bernal-stacked (AB) graphene exhibits the highest thermal conductivity. Introducing a twist reduces this property: Li et al. Li et al. (2014) measured a 25% decrease at $34^{\circ}$, while Han et al. Han et al. (2021) observed a 15% decrease at $2^{\circ}$, both using optothermal Raman spectroscopy. These reductions are attributed to changes in low-frequency phonon transport and are supported by molecular dynamics simulations. Other studies Nie et al. (2019); Wang, Xie, and Chen (2017); Li et al. (2018a) confirm that both in-plane and cross-plane thermal conductivities decrease as the twist angle increases.
@@
Here, we investigate the local thermal properties of twisted double bilayer graphene (TDBG) using Scanning Thermal Microscopy (SThM). We observe a reduction in thermal resistance of TDBG by $0.3\pm 0.1\times 10^{6}$ KW-1 compared to untwisted double bilayer graphene. This reduction is attributed to changes in both thermal conductivity and tip–sample thermal interface. Our findings are supported by analytical modeling.
Twisted double bilayer graphene consists of two AB-stacked bilayers placed on a 40 nm hexagonal boron nitride (hBN) substrate. The full structure is shown in Figure 1a. When stacked, a small twist angle $\theta$ ($<1^{\circ}$) is introduced between the bilayers, forming a moiré pattern of ABAB and ABCA domains, while suppressing the less stable ABBC regionsYoo et al. (2019); Li et al. (2021). These domains are separated by saddle point (SP) boundaries. The TDBG was fabricated using a dry transfer method with a polycarbonate (PC) film (10% concentration) supported by a polydimethylsiloxane bubble stamp. Two halves of a bilayer graphene, pre-cut using an atomic force microscopeLi et al. (2018b), were picked up sequentially with a relative twist angle of less than $1^{\circ}$. The final stack was transferred onto a hBN flake at room temperature without melting the PCGadelha et al. (2021), minimizing contamination during the process.
@@
Figure 3b shows a Scanning Thermal Microscopy image of the first twisted graphene region. The scan was taken across the boundary between the twisted and non-twisted areas to allow direct comparison of the thermal signal. As seen in the image, the thermal resistance increases in the twisted area by $\Delta R_{X}=0.3\pm 0.1\times 10^{6}$ KW-1. Topography and friction images reveal no visible differences between the two regions, confirming that the observed thermal contrast is not related to surface morphology or frictional properties.
A line profile extracted from the SThM image is shown in Figure 3c, along with a simultaneously acquired topography profile. A clear step in thermal resistance is visible at the twisted–non-twisted boundary, while no notable change is observed in the topography, aside from a slight graphene ripple marking the interface.
@@
The interface resistance ($R_{int}$) is often a dominant and uncertain factor in SThM measurements. It can be understood as the Kapitza resistance between the phonon systems of the probe and the sample. This resistance is influenced by material properties on both sides of the interface and any surface topography variations. In our case, the probe transitions from untwisted to twisted graphene. As shown in Figure 3c, no significant topography change is detected at this boundary.
@@
Finally, we consider the contribution of the sample’s thermal spreading resistance ($R_{spr}$) to the overall measured variation. Given the high thermal conductivity of hBN, we assume it is well thermalized with the underlying silicon oxide substrate. Therefore, the relevant thermal spreading path is defined by the graphene–on–hBN stack, as illustrated in Figure 3a. Analytical models for thermal spreading resistance in layered structuresEvangeli et al. (2019); Spiece et al. (2022) can be applied here. These models include several key parameters (see Supplementary Information for full details): the contact radius $a$, the in-plane ($k_{\parallel}$) and cross-plane ($k_{\perp}$) thermal conductivities of graphene, the graphene–hBN interface resistivity ($r_{int}^{gr-hBN}$), and the thermal conductivity of hBN ($k_{hBN}$).
Figure 3: (a) Thermal resistance map showing higher resistance on the twisted region compared to the non-twisted one. Scale bar is 500 nm. (b) Thermal resistance and topography profiles taken across the transition between the two regions.
Next, we consider the in-plane and cross-plane thermal conductivities of the twisted graphene layers. Several studiesWang, Huang, and Lu (2013); Zhang et al. (2017); Karak et al. (2021) have shown that graphene partially recovers its exceptional thermal transport properties when supported by hBN. The thermal conductivity of suspended single-layer graphene has been measured to reach up to 3000 Wm-1K-1Balandin et al. (2008); Pop, Varshney, and Roy (2012). When supported on high-quality boron nitride, reported values approach this suspended limit, typically ranging between 1000 and 2000 Wm-1K-1Pop, Varshney, and Roy (2012). For the cross-plane direction, bulk graphite exhibits a thermal conductivity of about 5 Wm-1K-1Pop, Varshney, and Roy (2012).
@@ Acknowledgements.
The authors acknowledge the following researchers for their support in the sample fabrication and measurements: S. Gonzalez-Munoz, K. Agarwal, P. de Crombrugghe de Picquendaele, Y. Hong, S. Mohapatra, K. Watanabe, T. Taniguchi, B. Nysten and B. Hackens. We acknowledge financial support from the F.R.S.-FNRS of Belgium (FNRS-CQ-1.C044.21-SMARD, FNRS-CDR-J.0068.21-SMARD, FNRS-MIS-F.4523.22-TopoBrain, FNRS-CR-1.B.463.22-MouleFrits, FNRS-PDR-T.0029.22-Moiré), from the Federation Wallonie-Bruxelles through the ARC Grant No. 21/26-116 and from the EU (ERC-StG-10104144-MOUNTAIN). This project (40007563-CONNECT) has received funding from the FWO and F.R.S.-FNRS under the Excellence of Science (EOS) programme. This work was also partly supported by the FLAG-ERA Grant TATTOOS, through F.R.S.-FNRS PINT-MULTI Grant No. R 8010.19. We acknowledge the technical support from Bruker UK. The support of the European Union’s Horizon 2020 Research Project and Innovation Program–Graphene Flagship Core3 (No. 881603), EPSRC EP/V00767X/1 HiWiN project is fully appreciated. R.R.-P. acknowledge the ERC starting grant TWISTRONICS. This work was supported by the French RENATECH network and the DIM-SIRTEC. K.W. and T.T. acknowledge support from JSPS KAKENHI (Grant Numbers 19H05790, 20H00354 and 21H05233).

### v2

@@ Abstract
Two-dimensional (2D) materials have attracted significant interest due to their tunable physical properties when stacked into homo- and hetero-structures. Twisting adjacent layers introduces moiré patterns that strongly influence the material electronic and thermal behavior. In twisted graphene systems, the twist angle critically alters phonon transport, leading to reduced thermal conductivity compared to Bernal-stacked configurations. However, experimental investigations into thermal transport in twisted structures remain limited. Here, we study the local thermal properties of twisted double bilayer graphene (TDBG) using Scanning Thermal Microscopy (SThM). We find an increase in thermal resistance of $0.3\pm 0.1\times 10^{6}$ K W-1 compared to untwisted bilayers, attributed to changes in both intrinsic thermal conductivity and the tip–sample interface. These results, supported by analytical modeling, provide new insight into thermal transport mechanisms in twisted 2D systems and offer a pathway toward thermal engineering in twistronic devices.
In recent years, two-dimensional (2D) materials have gained increasing attention. The atomic layers in these materials are held together by weak van der Waals forces, making them easy to cleave and stack into new heterostructuresGeim and Grigorieva (2013). Stacking different 2D layers can result in artificial systems with novel physical propertiesSutter and Sutter (2021); Li, Zhou, and Zhang (2017); Wu et al. (2022). Moreover, by twisting the layers with respect to each other, the periodicity of the resulting lattice changes, forming a so-called moiré patternHe et al. (2021). This twist angle becomes a crucial parameter for tuning the band structureRen et al. (2020) and the lattice symmetry of the new materialCarr et al. (2017); Z. and S. (2021). At small twist anglesKerelsky et al. (2019), the mismatch between layers leads to atomic reconstruction, favoring energetically stable stacking configurations. This results in discrete stacking domains and domain wallsYoo et al. (2019), which are predicted to significantly affect the local thermal properties of twisted 2D systems.
The thermal conductivity of graphene is known to depend on the twist angle between its layers. Bernal-stacked (AB) graphene exhibits the highest thermal conductivity. Introducing a twist reduces this property: Li et al. Li et al. (2014) measured a 25% decrease at $34^{\circ}$, while Han et al. Han et al. (2021) observed a 15% decrease at $2^{\circ}$, both using optothermal Raman spectroscopy. These reductions are attributed to changes in low-frequency phonon transport and are supported by molecular dynamics simulations. Other studies Nie et al. (2019); Wang, Xie, and Chen (2017); Li et al. (2018a); Krisna, Kawakami, and Koshino (2025) confirm that both in-plane and cross-plane thermal conductivities decrease as the twist angle increases.
@@
Here, we investigate the local thermal properties of twisted double bilayer graphene (TDBG) using Scanning Thermal Microscopy (SThM). We observe an increase in the thermal resistance of TDBG by $0.3\pm 0.1\times 10^{6}$ KW-1 compared to untwisted double bilayer graphene. This increase is attributed to changes in both thermal conductivity and tip–sample thermal interface. Our findings are supported by analytical modeling.
Twisted double bilayer graphene consists of two AB-stacked bilayers placed on a 40 nm hexagonal boron nitride (hBN) substrate. A scheme of the full structure is shown in Figure 1a. When stacked, a small twist angle $\theta$ ($1^{\circ}$) is introduced between the bilayers, forming a moiré pattern of ABAB and ABCA domains, while suppressing the less stable ABBC regionsYoo et al. (2019); Li et al. (2021). The TDBG was fabricated using a dry transfer method with a polycarbonate (PC) film (10% concentration) supported by a polydimethylsiloxane bubble stamp. Two halves of a bilayer graphene, pre-cut using an atomic force microscopeLi et al. (2018b), were picked up sequentially with a relative twist angle of less than $1^{\circ}$. The final stack was transferred onto a hBN flake at room temperature without melting the PCGadelha et al. (2021), minimizing contamination during the process.
@@
Figure 3a shows a Scanning Thermal Microscopy image of the first twisted graphene region. The scan was taken across the boundary between the twisted and non-twisted areas to allow direct comparison of the thermal signal. As seen in the image, the thermal resistance increases in the twisted area by $\Delta R_{X}=0.3\pm 0.1\times 10^{6}$ KW-1. Topography and friction images reveal no visible differences between the two regions, confirming that the observed thermal contrast is not related to surface morphology or frictional properties.
A line profile extracted from the SThM image is shown in Figure 3b, along with a simultaneously acquired topography profile. A clear step in thermal resistance is visible at the twisted–non-twisted boundary, while no notable change is observed in the topography, aside from a slight graphene ripple marking the interface.
@@
The interface resistance ($R_{int}$) is often a dominant and uncertain factor in SThM measurements. It can be understood as the Kapitza resistance between the phonon systems of the probe and the sample. This resistance is influenced by material properties on both sides of the interface and any surface topography variations. In our case, the probe transitions from untwisted to twisted graphene. As shown in Figure 3b, no significant topography change is detected at this boundary.
@@
Finally, we consider the contribution of the sample thermal spreading resistance ($R_{spr}$) to the overall measured variation. Given the high thermal conductivity of hBN, we assume it is well thermalized with the underlying silicon oxide substrate. Therefore, the relevant thermal spreading path is defined by the graphene–on–hBN stack, as illustrated in Figure 2a. Analytical models for thermal spreading resistance in layered structuresEvangeli et al. (2019); Spiece et al. (2022) can be applied here. These models include several key parameters (see Supplementary Information for full details): the contact radius $a$, the in-plane ($k_{\parallel}$) and cross-plane ($k_{\perp}$) thermal conductivities of graphene, the graphene–hBN interface resistivity ($r_{int}^{gr-hBN}$), and the thermal conductivity of hBN ($k_{hBN}$).
[…]
Next, we consider the in-plane and cross-plane thermal conductivities of the twisted graphene layers. Several studiesWang, Huang, and Lu (2013); Zhang et al. (2017); Karak et al. (2021) have shown that graphene partially recovers its exceptional thermal transport properties when supported by hBN. The thermal conductivity of suspended single-layer graphene has been measuredBalandin et al. (2008); Pop, Varshney, and Roy (2012) to reach up to 3000 Wm-1K-1. When supported on high-quality boron nitride, reported valuesPop, Varshney, and Roy (2012) approach this suspended limit, typically ranging between 1000 and 2000 Wm-1K-1. For the cross-plane direction, bulk graphite exhibitsPop, Varshney, and Roy (2012) a thermal conductivity of about 5 Wm-1K-1.
@@ Acknowledgements.
We acknowledge financial support from the F.R.S.-FNRS of Belgium (FNRS-CQ-1.C044.21-SMARD, FNRS-CDR-J.0068.21-SMARD, FNRS-MIS-F.4523.22-TopoBrain, FNRS-CR-1.B.463.22-MouleFrits, FNRS-PDR-T.0029.22-Moiré), from the Federation Wallonie-Bruxelles through the ARC Grant No. 21/26-116 and from the EU (ERC-StG-10104144-MOUNTAIN). This project (40007563-CONNECT) has received funding from the FWO and F.R.S.-FNRS under the Excellence of Science (EOS) programme. This work was also partly supported by the FLAG-ERA Grant TATTOOS, through F.R.S.-FNRS PINT-MULTI Grant No. R 8010.19. We acknowledge the technical support from Bruker UK. The support of the European Union’s Horizon 2020 Research Project and Innovation Program–Graphene Flagship Core3 (No. 881603), EPSRC EP/V00767X/1 HiWiN project is fully appreciated. R.R.-P. acknowledges the Flag-Era JTC project TATTOOS (N R.8010.19) and the Pathfinder project “FLATS” N 101099139. Sample fabrication was done within the C2N micro nanotechnologies platforms and partly supported by the RENATECH network and the General Council of Essonne. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) , the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan.
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.